OpenStudy (anonymous):

Conic Sections. -39+25x^2+50x+4y^2-24y=0

OpenStudy (arhin):

owusu, are u there

OpenStudy (anonymous):

[(5x)^2+2(5x)(5)+5^2]+[((2y)^2-2(2y)(6)+6^2]=39+25+36=100 (5x+5)^2+(2y-6)^2=100 Center(-5,6) Radius=square root of 100=10